Calculate the quantum of energy in (a) yellow light wavelength 6×10^_7 m (b) X rays of wavelength 5×10^10m
1
Expert's answer
2020-04-13T10:04:37-0400

 The quantum of energy is "E=h\nu" where "\nu=\frac{c}{\lambda}" and "h=6.6\u22c510^{\u221234}J\u00b7s" .


(a) For yellow light "\lambda_y=6\cdot 10^{-7}m" and "\nu_y=\frac{3\cdot10^8ms^{-1}}{6\cdot10^{-7}m}=5\cdot10^{14}Hz".


Quantum energy is "E_y=6.6\cdot 10^{-34}\cdot 5\cdot 10^{14}J=3.3\cdot 10^{-19}J"


(b) For X-ray we have "\lambda_x=5\cdot 10^{-10}m" , "\nu_x=\frac{3\cdot10^8ms^{-1}}{5\cdot10^{-10}m}=6\cdot10^{17}Hz" and


"E_x=6.6\cdot 10^{-34}\cdot 6\cdot 10^{17}J=3.96\cdot 10^{-16}J"


Answer:  the quantum of energy in (a) yellow light is "3.3\cdot 10^{-19}J" (b) X rays is "3.96\cdot 10^{-16}J"